Please help me to answer this problem.
"Two angles are complementary and the measure of one angle is 40degrees less than the measure of the other. How many degrees are in each angle?"
THANK YOU!

Answer by nerdybill(7384)

"Two angles are complementary and the measure of one angle is 40degrees less than the measure of the other. How many degrees are in each angle?"
.
Two angles are complementary if the sum is 90 degrees.
.
Let x = one of the two angles
then
x-40 = other angle
.
x + x-40 = 90
x + x-40 = 90
2x = 130
x = 65 degrees
.
Other angle:
x-40 = 65-40 = 25 degrees